# A broken debris cascade as a possible source of hot dust emission in transitioning planet-forming disks

### v1

@@ 5.2 Timescales
Figure 4: The settling (dotted lines), damping (dashed lines) and collisional (solid lines) timescales for different gas densities. The size below which damping becomes quicker than collisions becomes smaller with decreasing gas density. Small particles below $s_{\mathrm{b}}$ (thin vertical lines) do not collide since they move with the gas, the settling which occurs instead is much slower.
In Fig. 4 we show the timescales at which particles settle or damp into the midplane and at which they collide if they were in the collisional cascade. As described in Sec. 2, small particles below $s_{\mathrm{b}}$ (given by the vertical lines) move with the gas, and are thus removed from the cascade. In this figure we can also see how intermediate size particles ($s_{\mathrm{b}}<s<s_{\mathrm{cc}}$), which can collide, are essentially removed from the cascade as well because they are damped into the midplane more quickly than they collide.
What might not be obvious is why the collisional timescales change with gas density. Since we require an optical depth of 1, a certain amount of small particles is required. For lower gas densities, $s_{\mathrm{cc}}$ increases, causing relatively smaller particles to be included in the production of small particles. This is more efficient than when only very large particles are producing small particles, therefore $f_{0,\mathrm{cc}}$ can be lower for the same production rate, and collisional timescales increase.
5.3 Influence of fragment size distribution
Figure 5: The required mass flux $\dot{M}_{\mathrm{in}}$ for an opacity $\tau_{\mathrm{g}}$ of 1 as a function of the gas surface density $\Sigma_{\mathrm{gas}}$ for different values of the fragment size distribution power $\gamma_{\mathrm{f}}$.
All the results shown above have the underlying assumption that the fragment size distribution follows a power-law with $\gamma_{\mathrm{f}}=-3$. As described in Sec. 4.4, this value can be easily modified by working out the integrals numerically instead of algebraically. This value not only influences the required mass flux, but affects how this mass flux changes with gas density as well. This is shown in Fig. 5, where we plot the required mass flux against the gas density for different values of $\gamma_{\mathrm{f}}$. As expected, lower values of $\gamma_{\mathrm{f}}$ are more efficient at the production of small particles, reducing the required mass flux significantly.
@@ 5.4 Caveats
In deriving our results, some crude assumptions have been made. We assumed the gas to be homogeneous in density, whereas in reality a vertically isothermal disk will have a Gaussian density profile around the midplane (Shakura & Sunyaev, 1973). This would cause particles which are just large enough to not bind to the gas high up in the cloud to do so when crossing the midplane. Therefore, these particles are removed from the higher regions of the cloud within a single orbital period. However, in the current picture, these particles would fall in the intermediate-size range, meaning they are damped into the midplane quicker than they collide and did not contribute to the production of smaller particles nor to the cloud’s optical depth in the first place. Furthermore, for low values of $\gamma_{\mathrm{f}}$, the largest particles in the collisional cascade dominate the production of small particles, making these intermediate-size particles even more irrelevant.
@@
Furthermore, we have ignored the presence of particles which have reached the midplane, while these might still act as collision partners for larger particles which are still on their inclined orbits around the central star. This would increase the chance of particles colliding in the midplane relative to higher up in the cloud, which would lead to an increased concentration of collisional fragments in the midplane. Since we require a large part of the star to be covered by dust grains, this would be detrimental to the efficiency of this system. However, most particles which end up in the midplane are small (through settling) or intermediate-size (through damping) and since only particles larger than $\epsilon s$ are capable of destroying a particle of size $s$, most larger particles which are still on their inclined orbits will not be affected by an encounter with these particles.
@@ 6 Conclusions
We have studied the effect collisions between parent bodies on strongly inclined orbits on the abundance of small grains in planet-forming disks that are in transition and show a decrease of the gas surface density along with significant near-IR excesses. We anchored our calculations by requiring the optical depth of the collisionally produced dust to be of order 1. Our findings are as follows:
@@
    In deriving these results, we assumed a fragmentation power-law size distribution with $\gamma_{\mathrm{f}}=-3$. This is a very conservative value, since values down to $-5$ or $-6$ are plausible for the collisional velocities we are dealing with, which would be more efficient at producing small particles. Lower values for this power-law decrease the required mass flux drastically by several orders of magnitude, down to values as low as $\sim 10^{-9}$ $M_{\oplus}$/yr. Such mass fluxes seem easily sustainable for extended periods of time in planet-forming disks.
@@
In summary, we have shown that a collisional cascade in a planet-forming disk with reduced but non-zero gas surface densities can produce observable amounts of dust for extended periods of time, making this mechanism a candidate for explaining warm infrared excess in transition disks.
@@ Acknowledgements.
Appendix A Time-dependence of the size distribution
To solve Eq. (22) for the size distribution of tightly coupled particles, we can substitute
@@ Appendix A Time-dependence of the size distribution
to rewrite the differential equation to
@@
This can be solved by integration, resulting in
@@
Now, in the limit where $t\gg t_{\textrm{settle}}$, the steady-state size distribution is
@@
Appendix B Dust settling & inclination damping
A particle above the midplane, suspended in a gas, experiences drag. Depending on the situation, the particle moves in either of two possible ways: (1) the particle is large enough to move through the gas and follow its inclined orbit around the central star. Over a multitude of orbits, drag causes the inclination to decline/dampen. (2) the particle couples to the gas, which itself moves parallel to the midplane, and starts to move with it. Gravity slowly pulls the particle downwards, causing it to settle into the midplane.
Only considering the forces in the (vertical) $z$ direction, the acceleration resulting from drag is given by dividing the Epstein drag force (Eq. 1) over the particle mass (Eq. 4), resulting in
$\frac{F_{\textrm{D}}}{m}$ $=-\frac{\frac{4\pi}{3}\rho_{\textrm{gas}}s^{2}v_{\textrm{th}}v}{\frac{4\pi}{3}\rho_{\bullet}s^{3}}$ (62)
@@ Appendix B Dust settling & inclination damping
Furthermore, as described by Armitage (2010), in the limit where $z\ll R$, the gravitational acceleration in $z$ direction is given by
@@
Using this, the particle’s height above the midplane $z$ is given by
@@
$r$ $=\frac{-t_{\textrm{fric}}^{-1}\pm\sqrt{t_{\textrm{fric}}^{-2}-4\Omega_{\textrm{K}}^{2}}}{2}$ (68)
@@
We can now distinguish the two different cases by the sign of the square root’s argument. The argument is negative if $t_{\textrm{fric}}\Omega_{\textrm{K}}>1/2$, which corresponds closely to the requirement for a particle not being bound to the gas given by Armitage (2010) (which uses $t_{\mathrm{fric}}\Omega_{\mathrm{K}}>1$). In this case an oscillating motion occurs in the $z$-direction: the particle follows it’s inclined orbit and periodically crosses the midplane. The angular frequency of this motion is $\sqrt{\Omega_{\mathrm{K}}^{2}-(2t_{\mathrm{fric}})^{-2}}$, meaning the particle’s orbit is slightly slowed down relative to the Keplerian orbital frequency $\Omega_{\mathrm{K}}$. The amplitude of this oscillation, corresponding to the orbit’s inclination, is exponentially decreases on a timescale
@@
In other words: a particle which is not tightly coupled to the gas is damped into the midplane on a timescale twice the friction timescale.
The other case, where $t_{\textrm{fric}}\Omega_{\textrm{K}}<1/2$, describes the settling of a particle into the midplane: the particle has left it’s inclined orbit, and instead moves with the gas while slowly approaching the midplane. The timescales at which this happens, given by $-1/r$ from Eq. (68), are
@@
In the case where $t_{\textrm{fric}}\Omega_{\textrm{K}}\ll 1/2$ (corresponding to a tight coupling to the gas), this evaluates to
@@
Since the complete solution is a linear combination of these two exponential functions, the slowest exponential decay will dominate on longer timescales. For $t_{\mathrm{fric}}\Omega_{\mathrm{K}}<1/\sqrt{2}$ the second is the largest, so settling occurs on a timescale

### v3

@@ 5.2 Timescales
Figure 4: Settling (dotted lines), damping (dashed lines), and collisional (solid lines) timescales for different gas densities. The size below which damping becomes quicker than collisions becomes smaller with decreasing gas density. Small particles below $s_{\mathrm{b}}$ (thin vertical lines) do not collide since they move with the gas, the settling which occurs instead is much slower.
In Fig. 4, we show the timescales at which particles settle or damp into the midplane and at which they collide if they were in the collisional cascade. As described in Sect. 2, small particles below $s_{\mathrm{b}}$ (given by the vertical lines) move with the gas and are thus removed from the cascade. In this figure we can also see how intermediate-size particles ($s_{\mathrm{b}}<s<s_{\mathrm{cc}}$), which can collide, are essentially removed from the cascade as well because they are damped into the midplane more quickly than they collide.
What might not be obvious is why the collisional timescales change with gas density. Since we require an optical depth of 1, a certain amount of small particles is required. For lower gas densities, $s_{\mathrm{cc}}$ increases, causing relatively smaller particles to be included in the production of small particles. This is more efficient than when only very large particles are producing small particles; therefore, the value of $f_{0,\mathrm{cc}}$ can be lower for the same production rate and the collisional timescales increase.
5.3 Influence of the fragment size distribution
Figure 5: Required mass flux, $\dot{M}_{\mathrm{in}}$, for an opacity, $\tau_{\mathrm{g}}$, of 1 as a function of the gas surface density, $\Sigma_{\mathrm{gas}}$, for different values of the fragment size distribution power, $\gamma_{\mathrm{f}}$.
All the results shown above have the underlying assumption that the fragment size distribution follows a power law with $\gamma_{\mathrm{f}}=-3$. As described in Sect. 4.4, this value can be easily modified by working out the integrals numerically instead of algebraically. This value not only influences the required mass flux, but affects how this mass flux changes with gas density as well. This is shown in Fig. 5, where we plot the required mass flux against the gas density for different values of $\gamma_{\mathrm{f}}$. As expected, lower values of $\gamma_{\mathrm{f}}$ are more efficient at the production of small particles, reducing the required mass flux significantly.
@@ 5.4 Caveats
In deriving our results, some crude assumptions have been made. We assumed the gas to be homogeneous in density, whereas in reality a vertically isothermal disk will have a Gaussian density profile around the midplane (Shakura & Sunyaev, 1973). This would cause particles which are just large enough to not bind to the gas high up in the cloud to do so when crossing the midplane. Therefore, these particles were removed from the higher regions of the cloud within a single orbital period. However, in the current picture, these particles would fall in the intermediate-size range, meaning they are damped into the midplane quicker than they collide and did not contribute to the production of smaller particles nor to the cloud’s optical depth in the first place. Furthermore, for low values of $\gamma_{\mathrm{f}}$, the largest particles in the collisional cascade dominate the production of small particles, making these intermediate-size particles even more irrelevant.
@@
Furthermore, we have ignored the presence of particles which have reached the midplane, while these might still act as collision partners for larger particles, which are still on their inclined orbits around the central star. This would increase the chance of particles colliding in the midplane relative to higher up in the cloud, which would lead to an increased concentration of collisional fragments in the midplane. Since we require a large part of the star to be covered by dust grains, this would be detrimental to the efficiency of this system. However, most particles which end up in the midplane are small (through settling) or intermediate-size (through damping) and since only particles larger than $\epsilon s$ are capable of destroying a particle of a size, $s$, most larger particles which are still on their inclined orbits will not be affected by an encounter with these particles.
@@ 6 Conclusions
We have studied the effect collisions between parent bodies on strongly inclined orbits on the abundance of small grains in planet-forming disks that are in transition and show a decrease in the gas surface density along with significant NIR excess. We anchored our calculations by requiring the optical depth of the collisionally produced dust to be of order 1. Our findings are as follows:
@@
    In deriving these results, we assumed a fragmentation power-law size distribution with $\gamma_{\mathrm{f}}=-3$. This is a very conservative value, since values down to $-5$ or $-6$ are plausible for the collisional velocities we are dealing with, which would be more efficient at producing small particles. Lower values for this power law lead to a decrease in the required mass flux drastically by several orders of magnitude, down to values as low as $\sim 10^{-9}$ $M_{\oplus}$/yr. Such mass fluxes appear to be easily sustainable for extended periods of time in planet-forming disks.
@@
In summary, we have shown that a collisional cascade in a planet-forming disk with reduced but non-zero gas surface densities can produce observable amounts of dust for extended periods of time. This would make this mechanism a prime candidate for explaining warm IR excess in transition disks.
@@ Acknowledgements.
Appendix A Time dependence of the size distribution
To solve Eq. (22) for the size distribution of tightly coupled particles, we can substitute:
@@ Appendix A Time dependence of the size distribution
to rewrite the differential equation as:
@@
This can be solved by integration, resulting in:
@@
Now, in the limit where $t\gg t_{\textrm{settle}}$, the steady-state size distribution is:
@@
Appendix B Dust settling and inclination damping
A particle above the midplane, suspended in a gas, experiences drag. Depending on the situation, the particle moves in either of two possible ways: (1) the particle is large enough to move through the gas and follow its inclined orbit around the central star. Over a multitude of orbits, drag causes the inclination to decline; (2) the particle couples to the gas, which itself moves parallel to the midplane, and starts to move with it. Gravity slowly pulls the particle downwards, causing it to settle into the midplane.
Only considering the forces in the (vertical) $z$ direction, the acceleration resulting from drag is given by dividing the Epstein drag force (Eq. 1) over the particle mass (Eq. 4), resulting in:
$\frac{F_{\textrm{D}}}{m}$ $=-\frac{\frac{4\pi}{3}\rho_{\textrm{gas}}s^{2}v_{\textrm{th}}v}{\frac{4\pi}{3}\rho_{\bullet}s^{3}},$ (62)
@@ Appendix B Dust settling and inclination damping
Furthermore, as described by Armitage (2010), in the limit where $z\ll R$, the gravitational acceleration in $z$ direction is given by:
@@
Using this, the particle’s height above the midplane, $z,$ is given by:
@@
$r$ $=\frac{-t_{\textrm{fric}}^{-1}\pm\sqrt{t_{\textrm{fric}}^{-2}-4\Omega_{\textrm{K}}^{2}}}{2},$ (68)
@@
We can now distinguish the two different cases by the sign of the square root’s argument. The argument is negative if $t_{\textrm{fric}}\Omega_{\textrm{K}}>1/2$, which corresponds closely to the requirement for a particle not being bound to the gas given by Armitage (2010) (which uses $t_{\mathrm{fric}}\Omega_{\mathrm{K}}>1$). In this case an oscillating motion occurs in the $z$-direction: the particle follows it’s inclined orbit and periodically crosses the midplane. The angular frequency of this motion is $\sqrt{\Omega_{\mathrm{K}}^{2}-(2t_{\mathrm{fric}})^{-2}}$, meaning the particle’s orbit is slightly slowed down relative to the Keplerian orbital frequency $\Omega_{\mathrm{K}}$. The amplitude of this oscillation, corresponding to the orbit’s inclination, is exponentially decreases on a timescale:
@@
In other words, a particle which is not tightly coupled to the gas is damped into the midplane on a timescale twice the friction timescale.
The other case, where $t_{\textrm{fric}}\Omega_{\textrm{K}}<1/2$, describes the settling of a particle into the midplane: the particle has left it’s inclined orbit, and instead moves with the gas while slowly approaching the midplane. The timescales at which this happens, given by $-1/r$ from Eq. (68), are:
@@
In the case where $t_{\textrm{fric}}\Omega_{\textrm{K}}\ll 1/2$ (corresponding to a tight coupling to the gas), this can be expressed as:
@@
Since the complete solution is a linear combination of these two exponential functions, the slowest exponential decay will dominate on longer timescales. For $t_{\mathrm{fric}}\Omega_{\mathrm{K}}<1/\sqrt{2}$, the second is the largest, so settling occurs on a timescale